OpenStudy (anonymous):

integrate of -9t*e^-t dt can u help plz

OpenStudy (anonymous):

Use integration by parts method!

OpenStudy (anonymous):

integral (f dg) = f g- integral (g df), where f = t, dg = e^(-t) dt, df = dt, g = -e^(-t):

OpenStudy (anonymous):

ok then ?

OpenStudy (anonymous):

All you need to do is plug the values of f and g into their relation given, which gives a function and another integral to calculate.

OpenStudy (anonymous):

ok let me try it but what u get for the answer

OpenStudy (anonymous):

= 9 e^(-t) t-9 integral e^(-t) dt

OpenStudy (anonymous):

i dont remember the equation of part integration

OpenStudy (anonymous):

I hope you know how to integrate e^(-t) dt :)

OpenStudy (anonymous):

ya is just -e^(-t)

OpenStudy (anonymous):

= 9 e^(-t) (t+1)+constant

OpenStudy (anonymous):

is that right ?

OpenStudy (anonymous):

i didnt get how u get 9 e^(-t) (t+1)+constant

OpenStudy (anonymous):

= 9 e^(-t) t+9 e^(-t)+constant = 9 e^(-t) (t+1)+constant